5. Draw a set of coordinate axes on a piece of graph paper. Label the x axis“Output” (0 to 20 units) and the y axis “Cost” (0 to 20 units). Plot the following (x, y) points on your graph: (0, 4); (2, 6); (4, 7); (7, 8); (9, 9); (11, 11); (13, 14). Connect these points with a smooth curve and label it “Total Cost.” Working from this curve, construct a total fixed cost curve and total variable cost curve for the same firm.

6. Draw a second set of coordinate axes on another piece of graph paper. Label the horizontal axis “Output” (0 to 20 units) and the vertical axis “Cost per Unit” (0 to 2 units, in tenths of a unit). Using as a basis the total cost, total variable cost, and total fixed cost curves you drew for Problem 5, construct the following curves on your new graph: marginal cost, average total cost, average variable cost, and average fixed cost.
Answer #1
Q C TFC TVC ATC AVC AFC MC
0 4 4 0
2 6 4 2 3.00 1.00 2.00 1.00
4 7 4 3 1.75 0.75 1.00 0.50
7 8 4 4 1.14 0.57 0.57 0.33
9 9 4 5 1.00 0.56 0.44 0.50
11 11 4 7 1.00 0.64 0.36 1.00
13 14 4 10 1.08 0.77 0.31 1.50

TFC = TC (when Q = 0)

TVC = C - TFC

ATC = C/Q , AFC = TFC/Q and AVC = TVC/Q


MC (nth unit) = (C (Qn) - C (Qn-1))/(Qn - Qn-1)